I need help on how to figure this out:
To determine the number of deer in a game preserve, a conservationist catches 804 deer, tags them and lets them loose. Later, 549 deer are caught; 183 of them are tagged. How many deer are in the preserve? Please help

Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

804/X=183/549
183X=804*549
183X=441,396
X=441,396/183
X=2,412 IS THE ESTIMATED NUMBER OF THE DEER POPULATION.

Answer by stanbon(75887)

To determine the number of deer in a game preserve, a conservationist catches 804 deer, tags them and lets them loose. Later, 549 deer are caught; 183 of them are tagged. How many deer are in the preserve?
----------------
population/(tagged&released) = (# caught)/(# tagged in # caught)
--------------
population/804 = 549/183
pop = 804[549/183)
pop = 2412 deer
----------------------
Cheers,
Stan H.